# Integer Lattice Gas with a sampling collision operator for the fluctuating Navier-Stokes Equation

### v1

@@ V Computational Efficiency
This equation can only approximate a system with $\omega\neq 1$, as to approach $\omega=1$ would require an infinite number of collisions. Thus, we only compare our algorithm to Blommel and Wagner’s algorithm for $\omega\neq 1$.
In the case of Fig. 15 (a), we compare Blommel’s algorithm to the lattice Boltzmann, the sampling lattice gas, along with the maximum possible efficiency of the sampling lattice gas for various average densities, all initialized with a lattice size of $L=100$ and run for $\omega=0.5$. The starting average density is initialized via Eq. (32), for flat $\bar{N}(x,0)$ and $u(x,0)=0$. The lattice gas systems are initialized using a Poisson distribution around these $f_{i}^{eq}$ values given by Eq. (31). The sampling lattice gas systems are run for $t=10,000$ timesteps to allow the computational cost of initializing the lookup tables to become negligible.
All systems are then timed for 10,000 timesteps, and are transformed into MLUPS by taking,
$t_{MLUPS}=\frac{Lt}{1,000,000t_{s}},$ (86)
where $t_{s}$ is the time taken to run in seconds. The simulations were run on a Dell XPS laptop with an Intel I5 CPU, and a clock speed of 1.6 GHz.
For the maximal lattice gas, a full sampling lattice gas system was run, saving the sampled $\pi$-values and binomial random numbers (for $\omega\neq 1$) into arrays, before repeating the simulation and reading all random variables from those arrays. Because Blommel and Wagner’s algorithm takes far longer to run than the others, we ran a linear regression on a smaller sample of data to extrapolate the efficiency of the algorithm. The explicitly measured data points are displayed, along with the linear regression itself, and due to the linear scaling, a logarithmic scale is necessary on the y-axis for proper resolution.
The sampling lattice gas and lattice Boltzmann timings are on the same order of magnitude, though the lattice gas is approximately a factor of five slower. The maximum efficiency for the lattice gas algorithm, however, is nearly a factor of three faster than the efficiency of the lattice Boltzmann algorithm. This indicates that the sampling lattice gas algorithm has the potential to surpass the lattice Boltzmann algorithm in efficiency should the sampling algorithm for the $\pi$-distribution and binomial distribution be improved enough. All of the aforementioned algorithms are much more efficient than Blommel and Wagner’s algorithm.
Fig 15 (b) repeats the above test for all algorithms aside from the Blommel and Wagner algorithm for $\omega=1$. Due to the lack of additional binomial sampling needed for an $\omega=1$ system, the sampling lattice gas algorithm is more efficient, and thus only ends up approximately a factor of two worse than the lattice Boltzmann algorithm, and the maximal efficiency of the lattice gas is just over a factor of three better. Due to Fig. 15 (b) not needing to account for Blommel and Wagner’s algorithm, a linear scale is able to be used for the y-axis. Both graphs show that our sampling lattice gas algorithm is indeed competitive with the lattice Boltzmann algorithm, and may become superior in terms of direct efficiency should it be further optimized in the future.
@@ VI Outlook
Integer lattice gases are a promising simulation method for fluctuating systems, and they are also directly derivable from Molecular Dynamics simulations via the MDLG method Parsa and Wagner (2017). This is why we were motivated to look into them more closely. They also form a physical microsystem that has a Boltzmann limit that can lead to novel insights into developing unconditionally stable lattice Boltzmann methods. Lattice gas methods generally contain additional physical content like correlations, that cannot be captured with current lattice Boltzmann methods, as we saw in section III. This is why we intend to continue to develop integer lattice gas approaches.
There are several remaining open problems when considering the development of the sampling lattice gas algorithm. The most immediately practical of these is the extension to two and three dimensional systems. This would allow the full recovery of the capabilities of the Blommel and Wagner collision operator, and would also allow deeper investigations into the scaling effects of the spuriously conserved parity of the $n_{0}$-values. The main hurdle for this development is in finding the proper method for distributing momentum across multiple axes while maintaining conservation.
Over-relaxation is a further open question, and is one with some progress towards its development. As discussed in Section III, Strand and Wagner have developed a method for simulating over-relaxation in the diffusive case Strand and Wagner (2022). However, this method involves a momentum-flipping operator, and thus applying it to hydrodynamic systems is non-trivial.
Adding forcing to the collision operator is another direction of potential study, which would allow the simulation of non-ideal systems. This addition may utilize the coarse-graining of a molecular dynamics simulation, as was done by Parsa and Wagner previously for the ideal lattice gas, to analyze the form of the forcing lattice gas collision operator Parsa and Wagner (2017).
@@
Finally, further optimization of the sampling lattice gas algorithm is needed in order to truly match the efficiency of or outperform the lattice Boltzmann algorithm for hydrodynamic simulations. Though the lattice Boltzmann algorithm may have the advantage of removing the noise from simulations inherently, those systems that require natural fluctuations would find lattice gas algorithms to be a superior algorithm for practical simulation.
@@ Acknowledgements.
We would like to thank Matteo Lulli for discussing with us the potential influence of correlations on the collision operator, as well as ways to analyze this correlated state more thoroughly for future work. We would also like to thank Victor Ambrus for pointing out the isothermal SOD shock tube solution Negro et al. (2019).

### v2

@@ V.1 Computational Efficiency Results
In the $\omega=1$ limit, $C$ goes to infinity, and thus $\omega=1$ cannot perfectly be achieved by Blommel and Wagner’s algorithm.
In Fig. 15 (b), the sampling lattice gas simulates an average of approximately 29.2 MLUPS for densities greater than or equal to 600 average particles per lattice site, and an average of approximately 49.7 MLUPS for densities less than 600 average particles per lattice site. As Fig. 15 (b) represents an under-relaxed system, which requires an additional binomial sampling step at each lattice site for the lattice gas, it follows that the efficiency will be lower. To minimize this inefficiency, we used a binomial sampling method from the GNU Scientific Library (GSL) [42].
The under-relaxed sampling lattice gas is approximately 4.19 and 3.53 times slower for high and low densities respectively compared to the sampling lattice gas for $\omega=1$. The under-relaxed lattice Boltzmann too is slightly slower than its counterpart at $\omega_{LB}=1$, however this is only by a small margin, simulating an average of approximately 156.5 MLUPS. This represents an efficiency drop of only approximately 4%, and thus the lattice Boltzmann for an under-relaxed system has an efficiency that is approximately 5.35 and 3.14 times faster than that of the lattice gas under the same circumstances.
Fig. 15 (b) also shows the extent to which our algorithm improves on Blommel’s algorithm. The lattice gas in the under-relaxed case is much faster than Blommel’s algorithm, including for low densities. The data point for Blommel and Wagner’s algorithm with the highest efficiency is already approximately 3.99 times less efficient than the equivalent point for the sampling lattice gas, and the measured data point with the lowest efficiency is approximately 96.2 times less efficient. This trend then continues in the linear regression performed on the measured Blommel data, as the data point with the lowest efficiency of the linear regressed data is approximately 777.3 times less efficient than the equivalent point on the sampling lattice gas.
Finally, we wished to show what the potential maximum increase in efficiency due to the sampling of randomly distributed numbers might be for our sampling lattice gas algorithm. As we are not experts in the efficient random sampling of probability distributions, we find it likely that there are improvements to our algorithm that we could have made but missed. Thus, we pre-generated a set of properly distributed random numbers for each simulation by running a full simulation once, saving all randomly generated numbers for the local equilibrium ensemble and the binomial distribution used to simulate the particle fraction. we then read off those random numbers while timing a second simulation.
In Fig. 15 (a), the sampling lattice gas with pre-generated random numbers had a relatively consistent average for all densities of approximately 708 MLUPS, which is approximately 3.99 times more efficient than the low density lattice gas, and is approximately 4.31 times more efficient than the lattice Boltzmann algorithm. For Fig. 15 (b), the sampling lattice gas with pre-generated random numbers had a relatively consistent average for all densities of approximately 387.1 MLUPS, which is approximately 7.78 times more efficient than the low density lattice gas, and is still approximately 2.47 times more efficient than the lattice Boltzmann algorithm. These results show that there is still a great deal of potential in improving the sampling algorithms used by our lattice gas algorithm, including the potential of becoming even more efficient than the equivalent lattice Boltzmann algorithm.
V.2 Implementation of the Sampling Lattice Gas Algorithm
It remains to discuss the exact form of the sampling lattice gas algorithm used for the efficiency tests in the prior section. All our algorithms are written in C, and the full algorithm that we will be discussing here can be found on our GitHub page [43].
Because the local equilibrium ensemble out of which we wish to sample is not a well known distribution, using a library such as the GSL in C is non-trivial [42]. Though packages such as the GSL have the ability to sample from arbitrary distributions, the full distribution must be known in advance, which is not the case for our system. This is due to the fact that the $N$ and $J$ values that define the local $\pi$ equilibrium ensemble are not known until after the streaming step of the prior timestep. The ideal solution to this issue would be a library of full distributions pre-calculated for every combination of $N$ and $J$ values, however this method requires either a large amount of memory (scaling with $O(N)$) or to read the distribution from a file for each collision.
Thus, we decided to utilize a dynamic lookup table. This method is similar to what was done in our prior paper on the diffusive sampling lattice gas, only with the added need to manage the $J$ value as well as the $N$ value [21]. We first initialized the lookup table as a two-dimensional triangular array with 2N12N-1 entries in each sub-array, with a likely maximum amount of particles that would be at one lattice site given by three times the average system density. Should a state occur beyond this maximum density, the table would be rebuilt with a higher maximum density in the algorithm. As the system was running, the approximate center point of the cumulative equilibrium distribution ($\sum_{\pi=J}P^{0}(\pi;N,J)\approx 0.5$) of the local equilibrium ensemble was saved in the lookup table. This then allowed us to then navigate to any part of the local equilibrium ensemble by utilizing the recursive formula from Eq. (26) and its inverse. Because this dynamic lookup table adds entries on the fly, over the course of longer runs generating new table entries will become rare.
Algorithm 1 PiGenerate(N,J)
int pi=J
double Sum=1
double Log=0
while pi<<N do
[…]
  Sum=Sum+Log
  pi=pi+2
end while
pi=J
Log=-$\ln$(Sum)
[…]
while Sum<<0.5 do
[…]
  Sum=Sum+Log
  pi=pi+2
end while
Save pi as Lookup.PiHalf
Save Log as Lookup.LogHalf
Save Sum as Lookup.SumHalf
Generating this lookup table of saved points requires calculating the whole local equilibrium ensemble for $N$ and $J$, as we do not have an explicit formula for the cumulative distribution, and Eq. (26) is not normalized by default. Alg. 1 provides a pseudocode representation of the function that generates a lookup table entry for a particular $N$ and $J$ pair.
The natural logarithms are used in Alg. 1 due to the fact that for large values of $N$, the unlikelihood of the $\pi$ values at the ends of the distribution leads to floating point errors, something that using the logarithms of these values helps curb. The lookup table itself is, as was described previously, a triangular array with an element for each pair of $N$ and $J$ values. It is initialized as a C structure with three saved values: SumHalf (type double), LogHalf (type double), and PiHalf (type int). This is because to save the closest point of the cumulative local equilibrium ensemble to 0.5 (SumHalf) and be able to iterate up and down the distribution later with Eq. (26) and its inverse, we must know the $\pi$ value at which the SumHalf value occurs, as well as the value of $P^{0}$(PiHalf;N,J);N,J), which we save as its natural logarithm for the reason discussed previously (LogHalf).
Algorithm 2 PiLookup(N,J)
if No Lookup Table exists for this N and J pair then PiGenerate(N,J)
end if
Generate (double) $r$, a uniformly distributed random number between 0 and 1
double Sum=Lookup.SumHalf (read from Lookup table)
double Log=Lookup.LogHalf (read from lookup table)
int pi=Lookup.PiHalf (read from the lookup table)
if r>>Sum then
  while Sum>>r do
[…]
   pi=pi+2
  end while
else
  while Sum<<r do
   pi=pi-2
   Log=Log-$\ln((\mathrm{N}-\mathrm{pi})(\mathrm{N}-\mathrm{pi}-1)/(4(\mathrm{pi}+2)^{2}+\mathrm{J}^{2}$))
[…]
  end while
end if
return pi.
Alg. 1 is then called each time a new lookup table element is required. These entries are then used to sample a new value of $\pi$ for the $N$ and $J$ values provided. Alg. 2 shows how the saved lookup table values are used to efficiently sample a new value of $\pi$ out of our equilibrium ensemble, using a similar process of logarithmic adding to Alg. 1.
Algorithm 3 Single Lattice Site Iteration Routine
if $\omega<1$ then
  for all velocities $v_{i}$ do
   int $n^{\omega}_{i}$=GSLBinomial($n_{i}$$\omega$)
   int $n_{i}^{uncol}=n_{i}-n_{i}^{\omega}$
  end for
[…]
  if $J\geq 0$ then
   int pi=PiLookup(N,J)
   $n_{-1}^{\omega}$=(pi-J)/2
   $n_{0}^{\omega}$=N-pi
   $n_{1}^{\omega}$= (pi+J)/2
  else
   J=-J;
   int $\pi$=PiLookup(N,J)
   $n_{-1}^{\omega}$=(pi+J)/2
   $n_{0}^{\omega}$=N-pi
   $n_{1}^{\omega}$= (pi-J)/2
  end if
  for all velocities $v_{i}$ do
   $n_{i}=n_{i}^{\omega}+n_{i}^{uncol}$
  end for
else
[…]
  if $J\geq 0$ then
   int pi=PiLookup(N,J)
   $n_{-1}$=(pi-J)/2
   $n_{0}$=N-pi
   $n_{1}$= (pi+J)/2
  else
   J-̄J;
   int pi=PiLookup(N,J)
   $n_{-1}$=(pi+J)/2
   $n_{0}$=N-pi
   $n_{1}$= (pi-J)/2
  end if
end if
The particles are now in a redistributed state and are ready to be streamed once all lattice sites are redistributed.
Alg. 2 is used within our full iteration algorithm to sample a new value of $\pi$ out of our equilibrium ensemble. An example of this collision step for a single lattice site can be seen in Alg. 3. This collision algorithm assumes that the $n_{i}$ values are stored as an array of integers that has global scope, and accounts for both $\omega=1$ and $\omega<1$ systems. This is a practical implementation of the under-relaxation algorithm discussed in Sec. II A. $\omega$ values greater than one are treated identically to $\omega=1$, as this algorithm does not include overrelaxation.
The function GSLBinomial represents the function from the GSL package that returns a binomially sampled random number if given the number of trials (particles) $N$ and the probability of success (collision fraction) $\omega$ [42]. GSLBinomial is not the actual function name, and more implementation is needed to use the GSL sampling method than is included here, but we do not wish to over-complicate this section with a full technical description of its implementation. Please consult the documentation for the GSL for more thorough details. The $n_{i}^{uncol}$ values, then, represent the total number of particles of each density not selected for collision by the GSLBinomial function.
As the local equilibrium $\pi$ distribution is identical for positive and negative values of $J$, we account for negative values of $J$ in Alg. 3 by flipping the signs on the $J$ values in Eqs. (8) and (10).
Algs. 1, 2, and 3 combined represent the full process of collision for our lattice gas algorithm. Alg. 3 is then repeated over all lattice sites, and then the particles are streamed between lattice sites. Neither of these steps are novel, and thus will not be extrapolated upon further here.
@@ VI Outlook
Integer lattice gases are a promising simulation method for fluctuating systems, and they can also be directly derivable from Molecular Dynamics simulations via the MDLG method [27]. This is why we were motivated to look into them more closely. They also form a physical microsystem that has a Boltzmann limit that can lead to novel insights into developing unconditionally stable lattice Boltzmann methods. Lattice gas methods generally contain additional physical content like correlations, that cannot be captured with current lattice Boltzmann methods, as we saw in section III. Therefore integer lattice gases are an important bridge between molecular representations and their approximation through lattice Boltzmann methods.
There are several remaining open problems when considering the development of the sampling lattice gas algorithm. The most immediately practical of these is the extension to two and three dimensional systems. We are working on this problem and hope to publish a working prototype for this in the near future. We anticipate that this approach will fully recover all of the capabilities of the Blommel and Wagner collision operator.
The main hurdle for this development is the increase in the number of non-conserved moments in higher dimensions. In a D2Q9 system, for example, there are six non-conserved moments, while in a D3Q27 system there are 24 non-conserved moments. The local equilibrium ensemble for the system will, therefore, require us to find the proper distributions out of which to sample these non-conserved moments. We estimate that the extension to a D2Q9 system would be approximately 6 times less efficient for $\omega=1$ systems, and a factor of approximately 3 times less efficient for under-relaxed systems, due to the increased requirement for sampling specific random numbers.
We can also consider extending the velocity space of the system as a whole, which we estimate would have similar increased computational costs relative to the number of velocities added. The more significant hurdle to extending into multi-speed lattices would be identical to that of extending to higher dimensions: finding the proper local equilibrium ensemble out of which to sample.
Overrelaxation is a further open question, crucial for many possible practical applications of the method. We anticipate publishing a promising approach for this for the one dimensional model presented in this paper shortly.
Adding forcing to the collision operator is another important extension of our approach, which would allow the simulation of non-ideal systems. This addition may utilize the coarse-graining of a molecular dynamics simulation, as was done by Parsa and Wagner previously for the ideal lattice gas, to analyze the form of the forcing lattice gas collision operator [27].
@@
Finally, further optimization of the sampling lattice gas algorithm is needed to truly match the efficiency of or outperform the lattice Boltzmann algorithm for hydrodynamic simulations. Though the lattice Boltzmann algorithm may have the advantage of removing the noise from simulations inherently, those systems that require natural fluctuations would find lattice gas algorithms to be a superior algorithm for practical simulation.
@@ Acknowledgements.
We would like to thank Matteo Lulli for discussing with us the potential influence of correlations on the collision operator, as well as ways to analyze this correlated state more thoroughly for future work. We would also like to thank Victor Ambrus for pointing out the isothermal SOD shock tube solution [39].